OpenStudy (anonymous):

A basket contains 8 pieces of fruit: 3 apples, 3 oranges, and 2 bananas. Jonas takes a piece of fruit at random from the basket, and then Beth takes a piece at random. What is the probability that Jonas will get an orange and Beth will get an apple?

OpenStudy (lucaz):

P(Jonas | orange)= #oranges / #fruits = 3/16 P(Beth | apple)= #apples / #fruits = 3/(16-1) = 3/15 = 1/5